OpenStudy (anonymous):

(√34)/(5√14) Can someone explain to me how to solve this??

OpenStudy (anonymous):

10√(17*7)

OpenStudy (anonymous):

5 √(34 * 14) = 5 √(17*2*2*7) = 5*2 √(17*7) = 10 √(17*7)

OpenStudy (anonymous):

10 √119

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

r u dividing or multiplying

OpenStudy (anonymous):

??

OpenStudy (anonymous):

if it is dividing / then √34 / 5√14 = √(2*17) / 5*√(2*7) √17 / 5√7